OpenStudy (anonymous):

There are 10 children in a room. The ratio of boys to girls increased when another boy and another girl entered the room. What is the greatest number of boys that could originally have been in the room? A. 1 B. 4 C. 5 D. 9 E. 10

OpenStudy (ankit042):

x/10-x initial ratio 0<=x<10 x+1/10-x+1 final ratio x/10-x < x+1/11-x Taking only numerator x(11-x)-(x+1)(9-x) denominator will always be >0 (Hint check the range of x) 11x -x^2 -9x -9 +x^2 > 0 2x -11 > 0 x > 5.5

hero (hero):

\[\frac{4}{6} \approx 66\%\] \[\frac{5}{7} \approx 71\%\]

OpenStudy (ankit042):

Calculation mistake :( 11x -x^2 -8x -9 +x^2 > 0 4x -9> 0 x > 9/4 so x can start from 3

hero (hero):

It's better in this case to calculate using actual numbers.

hero (hero):

Clearly the initial ratio that works is \[\frac{\text{boys}}{\text{girls}} = \frac{4}{6}\]

hero (hero):

Because if you increase each by one you get \(\large\frac{5}{7}\) and \[\frac{5}{7} > \frac{4}{6}\]

OpenStudy (ankit042):

Is it only me or you guys are facing difficulty in viewing text written from equation tab? I am seeing \(\large\frac{5}{7}\) and \[\frac{5}{7} > \frac{4}{6}\] It is not parsed properly

OpenStudy (ybarrap):

\[B+G=10\] \[B=10-G\] \[\frac{ B+1 }{ G+1 }>\frac{ B }{ G }\] \[\frac{10-G+1}{G+1}> \frac{10-G}{G}\] \[G(11-G)>(10-G)(G+1)\] \[11G-G^{2}>9G-G^{2}+10\] \[11G>9G+10\] \[2G>10\] \[G>5\] So G=6,B=10-G=4 Ans. 4

OpenStudy (anonymous):

Thanks everyone .. But I'm a little confused...

hero (hero):

@ybarrap that's great if you want to do more work than you need to.

hero (hero):

In addition to confusing the user

hero (hero):

@ankit042 , I'm able to view everything just fine. What browser are you using?

OpenStudy (anonymous):

Thanks to everyone. I understand now :)

OpenStudy (ybarrap):

np

hero (hero):

Which explanation did you understand?

hero (hero):

I think that she just understood that the answer is 4.

OpenStudy (ankit042):

I am using chrome Version 28.0.1500.71 on ubuntu 12.04

hero (hero):

I'm using Chrome...same version so it must be the unbuntu.

OpenStudy (ankit042):

hmm maybe! It used to work before though!

OpenStudy (ybarrap):

Refresh the page. This may refresh the java libraries associated with LaTex

OpenStudy (anonymous):

I understood @ybarrap's explanation